OpenStudy (anonymous):

I have a vector field F(x,y) and a hyperbola; I'm supposed to calculate the work from the field on a particle traveling on the right curve of the hyperbola; how do you go about deciding a good parametrization?

OpenStudy (anonymous):

For example: 1/4*x^2-1/9*y^2=1 from (2*sqrt(5),-6) to (2*sqrt(2),3)

ganeshie8 (ganeshie8):

lets try... i did line integrals very long back.... is that a real example problem ? if so, where is the vector field ?

OpenStudy (anonymous):

The vector field is (1/4*x^2,1/6*x^2)

OpenStudy (anonymous):

Thoughts? :D

OpenStudy (anonymous):

isnt the work is the derevative ?

OpenStudy (anonymous):

\[\int\limits_{\alpha}^{\beta}(P(x(t),y(t))x'(t)+Q(x(t),y(t))y'(t))dt\] Is the work of the field F=(P,Q) along some curve (on differential form).

OpenStudy (anonymous):

sorry i dont know , still learning these stuff :)

ganeshie8 (ganeshie8):

a dumb way wud be to parameterize :- x = 2cosht y = 3sinht

OpenStudy (anonymous):

Why dumb?

ganeshie8 (ganeshie8):

uhmm lets try it

OpenStudy (anonymous):

Oh, you mean it's a hassle? :d

OpenStudy (anonymous):

x= 2 sec(t) y= 3 tan(t)

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

x=2cosh(t),y=3sinh(t) gives a crazy t interval, no? :c

ganeshie8 (ganeshie8):

logs take over :o

ganeshie8 (ganeshie8):

try sec and tan..

OpenStudy (anonymous):

What about the interval for t?

OpenStudy (anonymous):

This shouldn't be so hard. :(

ganeshie8 (ganeshie8):

yup! i think logs wud make hyperbolics easy... finish the hyperbolic parametrization first maybe...

OpenStudy (anonymous):

log(sqrt(5)+sqrt((sqrt(5)-1) (1+sqrt(5))))<=t<=log(sqrt(2)+sqrt((sqrt(2)-1) (1+sqrt(2)))) This does not feel right.

ganeshie8 (ganeshie8):

\(-\log (2+\sqrt{5}) \le t \le -\log (\sqrt{2} - 1) \)

OpenStudy (anonymous):

Supposedly not right. :/

OpenStudy (anonymous):

This problem is driving me nuts!

ganeshie8 (ganeshie8):

wat do u mean not right ?

ganeshie8 (ganeshie8):

it doesnt look that pleasant, but it does look right to meh

OpenStudy (anonymous):

Apparently not :C

OpenStudy (anonymous):

The answer is supposedly gnarly though.

ganeshie8 (ganeshie8):

hmm-

OpenStudy (anonymous):

Is your field \[ \left ( \frac {x^2}4, \frac {x^2}6 \right) \] There is no y?

OpenStudy (anonymous):

There is no y.

OpenStudy (anonymous):

If the second component is in terms of y, then it would be very easy. That is why I am asking.

OpenStudy (anonymous):

I think that your teacher should not ask such a question that yields to complicated integrals.

OpenStudy (anonymous):

I tend to agree, the problem is generated through RNG though.

OpenStudy (anonymous):

What is RNG?

OpenStudy (anonymous):

Random number generator

OpenStudy (anonymous):

With restrictions of course.

OpenStudy (anonymous):

There is for sure one trick to make this easy.

OpenStudy (anonymous):

It should be known, that between all integrals that you can generate, very few can be done by hand. The others might need numerical approximations.

OpenStudy (anonymous):

It's generated to be hand-solved with relative ease.

OpenStudy (anonymous):

The trick that I was looking for is was that the field is conservative, but it is not.

OpenStudy (anonymous):

If you find an easy solution, please let us know.

OpenStudy (anonymous):

Here is an idea \[ \left ( \frac {x^2}4, \frac {x^2}6 \right)= \left ( \frac {x^2}4, 0 \right)+\left ( 0, \frac {x^2}6 \right) \]

OpenStudy (anonymous):

The first one is conservative and the work for it depends of the end points. You need to complete the work done by the second.

OpenStudy (anonymous):

You need to compute \[ \int_{t_1}^{t_2} \frac 1 6 x(t)^2 y'(t) dt \]

OpenStudy (anonymous):

Try it